40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 243 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  4402   Thu Mar 10 17:03:48 2011 Larisa ThorneConfigurationElectronicscalculations for passive low pass filter on X arm

[Kiwamu, Larisa] 

 

We want to increase gain in the lower frequencies, so a circuit must be designed (a passive low pass filter). 

 

First, measurements were taken at the X arm for impedance and capacitance, which were 104.5kOhms and 84.7pF respectively. Kiwamu decided to make the circuit resemble a voltage divider for ease of calculation, such that Vout/Vin would be a ratio of some values of the equivalent circuit reactance values. After a few algebra mistakes, this Vout/Vin value was simplified in terms of the R, C measured and the R', C' that would be needed to complete the circuit. 

Since the measured C was very small and the measure R was fairly high, the simplified form allowed us to pick values of R' and C' that would make the critical frequency occur at 0.1Hz: set the R' resistance to 1MOhm and C' capacitance to 10uF, which would yield a gain ~1.

With these values a circuit we can start actually making the circuit.

  4401   Thu Mar 10 16:00:53 2011 Aidan, JoeUpdateGreen LockingIntensity stabilization loop using beatnote DC.

Aidan: Joe and I have added a channel that takes the DC output from the vertex beatnote PD and sends it, via RFM, to a DAC at the ETMX end. Immediately before the output is a filter C1GCX_AMP_CTRL. The output of the DAC is connected to the CURRENT LASER DIODE modulation input on the back of the Innolight driver. This will modulate the current by 0.1A/V input.

We should be able to modulate the green laser on the end now and stabilize the intensity of the amplitude on the beatnote PD at the vertex. (In this configuration, the ampltiude noise of the PSL laser will be injected onto the end laser - we may want to revisit that).

Joe's comments on model change:

I added a RFM connection at the output of the C1:GCV-XARM_BEAT_DC filter in the c1gcv model.  The RFM connection is called: C1:GCV-SCX_ETMX_AMP_CTRL.

This RFM connection goes to the c1scx model and into Kiwamu's GCX box, which uses top_names.  There's a filter inside called AMP_CTRL, so the full filter name is C1:GCX-AMP_CTRL.  The output then goes to the 7th DAC output.

Photos:

  1. NPRO CURRENT CTRL plugged into DAC channel 7
  2. You can actually see it's channel 7 in this image
  3. The other end plugged into the back of the Innolight driver
  4. Schematic of the setup
  5. Updated C1ALS_OVERVIEW MEDM screen (I don't know why the field in the background turned orange - maybe it's coming into a long dry summer?)

 

Quote:

There are 3 standard techniques to reduce this effect:

1) Stabilize the end laser by sensing the green light coming into the PSL before recombination and feeding back with SR560 (this is the only one that you should try at first).

 

The reason that I chose this PD is that, apparently, the green light coming from the cavity is clipped when it is picked off for its DC PD.

Attachment 1: P1000313.jpg
P1000313.jpg
Attachment 2: P1000314.jpg
P1000314.jpg
Attachment 3: P1000315.jpg
P1000315.jpg
Attachment 4: GREEN_ISS_LOOP.pdf
GREEN_ISS_LOOP.pdf
Attachment 5: Screenshot-C1ALS_OVERVIEW.adl.png
Screenshot-C1ALS_OVERVIEW.adl.png
  4400   Thu Mar 10 14:30:53 2011 ranaUpdateGreen LockingIntensity noise limits the beatnote sensitivity

There are 3 standard techniques to reduce this effect:

1) Stabilize the end laser by sensing the green light coming into the PSL before recombination and feeding back with SR560 (this is the only one that you should try at first).

2) Moving to the center of the MFD fringe via ETM steps.

3) Auto-alignment of the beam to the arm.

  4399   Thu Mar 10 14:29:05 2011 KojiUpdateGreen LockingIntensity noise limits the beatnote sensitivity

We can modify the freq divider circuit to make it a comparator.

Quote:

The next steps we should do are :

    - to investigate a cause of the intensity fluctuation
          * end green laser
          * suspensions' angular motions
          * doublecheck the RIN contribution if it's from the PSL or the X arm in the beatnote RFPD to make sure the RIN is dominated by the X arm transmitted light
  
    - to think about how to make the system insensitive to the intensity noise
          - bring the beat frequency to the zero cross point of the MFDs ?
          - PLL ?

Quote:

We are limited by the intensity noise of the X arm transmitted green light.

 

  4398   Thu Mar 10 14:22:58 2011 kiwamuUpdateGreen LockingIntensity noise limits the beatnote sensitivity

The next steps we should do are :

    - to investigate a cause of the intensity fluctuation
          * end green laser
          * suspensions' angular motions
          * doublecheck the RIN contribution if it's from the PSL or the X arm in the beatnote RFPD to make sure the RIN is dominated by the X arm transmitted light
  
    - to think about how to make the system insensitive to the intensity noise
          - bring the beat frequency to the zero cross point of the MFDs ?
          - PLL ?

Quote:

We are limited by the intensity noise of the X arm transmitted green light.

  4397   Thu Mar 10 14:06:54 2011 kiwamuUpdateGreen LockingIntensity noise limits the beatnote sensitivity

We are limited by the intensity noise of the X arm transmitted green light.

Since the intensity noise from the PSL wasn't big enough to explain the differential noise (#4392), so this time I measured the noise contribution from the X arm transmitted light.
diff_noise_Mar8.png

 


(coupling measurement)

  IN_TF_complete.png

  I performed the same intensity noise coupling measurement, but this time between the DC signal of the beatnote RFPD and the beatnote signal.

 While measuring it, I excited the intensity of the PSL laser by using the same AOM like I did yesterday. This AM cause the observable intensity noise on the beatnote RFPD.

With the excited AM, we can pretend to have an excited AM on the green transmitted light from the X arm, of course assuming the intensity noise coupling from the PSL is less.

  4396   Thu Mar 10 13:44:56 2011 josephbUpdateCDSAdded digitization noise to the c1spy model for simulated ADCs/DACs

To simulate digitization noise, the easiest way I found was to use the MathFunction block, found in the CDS_PARTS model, under simLinkParts. 

The MathFunction block supports square of input value, square root of input value, reciprocal of input value, and modulo of two input values.

The last is useful because it casts the input values as integers before taking the modulo.By placing this block after the saturation block (set to +/- 32768), adding 32768.5, choosing the 2nd input to be larger than 2 * 32768 (100,000 in this case), and then subtracting 32768, we wind up with a rounding function. 

The above method has been applied to the c1spy model in the CI and SO out sub-blocks.

  4395   Thu Mar 10 01:31:37 2011 KevinUpdateElectronicsAS55 Characterizations

I measured the transfer function, shot noise, and dark spectrum of AS55.

From the shot noise measurement, the RF transimpedance is (556.3 +- 0.8) Ohms and the dark current is (2.39 +- 0.01) mA. The dark noise agrees with the approximate value calculated from the circuit components.

There are no anomalous oscillations when there is no light on the photodiode. I am working on fitting the transfer function in LISO but the other plots are on the wiki at http://blue.ligo-wa.caltech.edu:8000/40m/Electronics/AS55

  4394   Thu Mar 10 01:28:47 2011 joe, jamie, rana, chrisSummaryCDSSimSuspension !

Today was a banner day for Simulated Plants.

Joe and Jamie have been working to get it all happening and this afternoon we started stuffing filters into the plant to make it act like the:

40mETMY.png

We put in the following features so far:

  1. Anti-Imaging filters (these are hacked to be approximate since the real ones are 7570 Hz LP filters and the SimAI only can have filters up to 8192 Hz).
  2. Dewhitening filters (copied from the SimDW in the SUS-ETMY screens)
  3. Coil Driver transimpedance (1 / 200 Ohms)
  4. Magnet-coil force constant (0.016 N/A)
  5. Conversion from Coil to DOF Basis
  6. All DOFs of the mechanical model are represented as simple harmonic oscillators with Q~100 and f ~ measured free swinging peaks.
  7. Signals/Noise can be injected either as force noise on the test mass or as displacement noise at the suspension point.
  8. Conversion from DOF to Shadow Sensor basis.
  9. Optical Levers (with whitening)
  10. Shadow Sensors have 2V/mm readout gain and whitening filters before being digitized by the SimADC.

We have also changed the switching logic for the SUS and SimETMs for the shadow sensor whitening. It used to be that either the hardware OR the software whitening was on. Now we have made it like all of the other whitening/antiwhitening in LIGO and the whitening/antiwhitening come on together. Joe and Jamie are going to propagate this to the other SUS. The hardware filter is a 30,100:3 (poles:zeros) whitening filter. The digital filter used to also be 30,100:3 with a DC gain = 1. I've changed the FM1 filter in the XXSEN filter banks into a 3:30 for the ETMY so that it now comes on and just compensates the hardware filter. This change should be propagated to all other SUS and the MEDM screens updated to show the new situation.

After this change, we decided to calibrate the {UL,UR,LL,LR,SD}SEN channels into units of microns. To do this we have made an FM6 filter called 'cts2um' that accounts for the OSEM gain and the ADC conversion factors. These channels are now in units of microns without applying any calibration in the DTT or Dataviewer. The plot below shows the OSEM shadow sensor time series with all damping loops ON and a very rough version of seismic noise being injected in all 6 DOFs (note that the y-axis is microns and the x-axis is seconds).

dvsim.png

Next, Jamie is adding the angular calibrations (so that SUSPIT and SUSYAW are in rads) and Chris is making vectift quality seismic noise injectors.

We also need to add coating thermal noise, suspension thermal noise, substrate thermal noise, ADC/DAC noise and a lot of MEDM screen indicators of what state we're in. I myself can't tell from the OSEM time series if its real or Sim.

redpill_bluepill.jpg

  4393   Wed Mar 9 23:19:04 2011 kiwamuUpdateCDSrebooted c1ioo

For some reason the c1ioo machine suddenly died just 30 miteus before.

It died after we added a DAQ channel for c1gcv and rebooted the frame builder.

It didn't respond to a ping command. Therefore I rebooted the machine by clicking the physical reset button.

Now it seems fine.

  4392   Wed Mar 9 18:17:11 2011 kiwamuUpdateGreen LockingIntensity noise coupling

Here is a new plot for the differential noise measurement. I plot a noise contribution from the intensity noise (brown curve).

If we believe this data, the differential noise is NOT dominated by the intensity noise of the PSL.

diff_noise.png

 


(intensity noise coupling measurement)

 Here is a plot for the transfer functions (TFs) from the intensity noise DCPD to the beat signal.

IN_TF.png

   In principle these TFs tell us how much intensity noise are contributed into the differential noise.

When I measured the spectra shown above, the frequency offset of the beatnote was at about 8 MHz from the zero cross point.

Keeping the same lock, I measured the transfer function (red curve) by using the swept sine technique on DTT. The setup for this measurement is depicted on the last entry (#4389).

Then I made the spectra above by multiplying the intensity spectrum by this TF.

  Later I measured another transfer function when the beatnote was at about 2 MHz from the zero cross point.

According to this measurement, our MFD gets insensitive to the intensity noise as the beat offset goes close to the zero cross point. This is consistent with what we expected.

  4391   Wed Mar 9 17:29:11 2011 steveSummaryVACsingle O-ring protection

We have one single O-ring on the 40m vacuum envelope. It is on the OOC west side, facing the AP table. This O-ring has to be protected from the force of this

door. There should be 3 shims  ~120 degrees apart to carry the full load, so it is not the O-ring that is getting squashed.

This morning I found only one of these shims in place.

Attachment 1: so1.jpg
so1.jpg
Attachment 2: P1070458.JPG
P1070458.JPG
  4390   Wed Mar 9 16:07:42 2011 kiwamuUpdateVIDEOcable session

[Koji, Steve, Suresh, Kiwamu]

The following video cables have been newly laid down :

  - MC1F/MC3F (65 ft.)

  - PMCR (100 ft.)

  - PSL spare (100 ft.)

  - PSL1  (100 ft.)

  - PSL2  (100 ft.)

 

  4389   Wed Mar 9 04:46:13 2011 kiwamuUpdateGreen Lockingmore intensity noise measurement

 

Here is a diagram for our intensity noise coupling measurement.

intensity_setup.png

 

The below is a plot for the intensity noise on the DCPD. (I forgot to take a spectra of the PD dark noise)

For some reason, the RIN spectrum becomes sometimes noisier and sometimes quieter. Note that after 10 pm it's been in the quiet state for most of the time.

An interesting thing is that the structure below 3 Hz looks like excited by motion of the MC when it's in the louder state.

IntensityNoise.png

Quote: from #4383

A photo diode and an AOM driver have been newly setup on the PSL table to measure the intensity noise coupling to the beat note signal.

We tried taking a transfer function from the PD to the beat, but the SNR wasn't sufficient on the PD. So we didn't get reasonable data.

  4388   Tue Mar 8 16:59:47 2011 josephbUpdateCDSSimulated Plant Work

The screens for the simplified c1spx model have been updated.  I re-introduced the suspension point information into the sensor output matrix so we can take into account the fact that as the entire supporting structure moves, the osems moves relative to the optic.

Master screens for the noise filters (i.e. 60 Hz, suspension point motion, and optic noise) have been created.

I have currently set the matrix values of the c1spx model to handle just longitudinal motion.  I.e. Coils drive only in the POS degree of freedom and sensor read outs are also only in the POS degree of freedom.  I've turned off all the noise inputs.

I added a simple double pole at 1 Hz in the C1:SUP_ETMX_PL_F2P_0_0 filter bank.

  4387   Tue Mar 8 15:33:09 2011 kiwamuSummaryGreen Lockingplan on Mar.8th
Today's goal is to measure the contribution from the intensity noise to the beatnote.
 
Plans for today
  - check the ADC for the DCPD that Jenne installed yesterday
  - adjust RF power on the AOM
  - take spectrum of the differential noise and measure the coupling from the intensity noise
  - update the noise budget

Quote: from #4382
This week's goal is to investigate the source of the differential noise and to lower it.

 

  4386   Tue Mar 8 15:23:16 2011 steveUpdatePEMclean room gloves

Ansell AccuTech 91-300 clean room gloves  ONLY in the 40m lab.

Cleaning and preparation must be carried out in these gloves also.

  4385   Tue Mar 8 15:20:31 2011 kiwamuUpdateGreen Lockingdifferential noise on Mar.8th

differential_noise20110308.png

Noise below 10 Hz became larger again compared with the data before (see here #4352)

Note that the Y-axis is in MHz.

  4384   Tue Mar 8 14:50:19 2011 kiwamuUpdateCDSnames for filter modules

[Joe/Kiwamu]

 We found there are some filter names that we can not properly build for some reason.

The following names are not properly going to be built :

 - REFL_DC

 - AUX

If we use the names shown above for filters, it doesn't compile any filter modules.

We took a quick look around the src files including feCodegen.pl, but didn't find any obvious bugs.

  4383   Tue Mar 8 06:29:06 2011 kiwamuUpdateGreen LockingIntensity noise setup

[Jenne, Chris, Kiwamu]

 A photo diode and an AOM driver have been newly setup on the PSL table to measure the intensity noise coupling to the beat note signal.

We tried taking a transfer function from the PD to the beat, but the SNR wasn't sufficient on the PD. So we didn't get reasonable data.

 

(what we did)

  - put a DCPD after the doubling crystal on the PSL table. The PD is sitting after the Y1 mirror, which has been used for picking off the undesired IR beam.

  - installed the AOM driver (the AOM itself had been already in place)

  - injected some signals onto the AOM to see if we can see an intensity fluctuation on the PD as well as the beat signal

 

(intensity noise)

  In order to have better SNR for the intensity measurement, we put an AC coupled SR560 with the gain of 100 just before the ADCs.

When a single frequency signal was applied from a Stanford Research's function generator to the AOM, we could clearly see a peak at the doubled frequency of the injected signal.

Also a peak at the same frequency was found on the beat note signal as well.

But when random noise was injected from the same function generator, the random noise looked below the ADC noise.

Jenne adjusted the output voltage from the PD to about 1 V to avoid a saturation in the analog path, but later we realized that the ADC counts was marely ~ 20 counts.

So we will check the ADC tomorrow. Hopefully we will get a good SNR.

  4382   Mon Mar 7 18:20:01 2011 kiwamuSummaryGreen Lockingplans
This week's goal is to investigate the source of the differential noise and to lower it.
 
Plans for tonight
 - realign GREEN_TRANS PD at the PSL table
 - update the noise budget
 - take spectrum of the differential noise
 - investigate a noise coupling to the differential noise especially from the intensity noise
 - update the noise budget again
 
Plans for this week :
 - Auto alignment scripts for green (Kiwamu)
 - connect the end REFL_DC  to an ADC (Kiwamu)
 - make an active phase rotation circuit for the end PDH (undergrads)
 - bounce-roll notches (Suresh)
 - optimization of the suspensions including the input matrices and the Q-values (Jenne)
 - optimization of MFSS (Koji/Rana/Larisa)
 - rewire the mechanical shutter on the 1X9 binary outputs (Steve)

 

  4381   Mon Mar 7 17:58:14 2011 sureshSummaryGeneralStuff from LLO

Here is the updated list. These lists were used as packing lists and therefore are organised by Box #.
Attachment 1: eLIGO_items_from_LLO_for_Caltech_Sheet1.pdf
eLIGO_items_from_LLO_for_Caltech_Sheet1.pdf eLIGO_items_from_LLO_for_Caltech_Sheet1.pdf eLIGO_items_from_LLO_for_Caltech_Sheet1.pdf eLIGO_items_from_LLO_for_Caltech_Sheet1.pdf eLIGO_items_from_LLO_for_Caltech_Sheet1.pdf eLIGO_items_from_LLO_for_Caltech_Sheet1.pdf
  4380   Mon Mar 7 17:22:39 2011 josephbUpdateCDSNew simulated plant work

[Joe, Jamie]

We modified the c1scx model to have a switch to go between simulated and real plants.  The channel is currently C1:SCX-SIM_SWITCH. 

When this channel is zero, the simulated plant channels are going to the ADCs and zeros  are going out to the real DACs.  When this channel is one, the real ADCs are coming in, and real data is going out to the DACs.

Jamie will be adding a big green/red light to the suspension screens which indicate the state of the simulated plant.  We will also eventually add this to the overall status screen.

A control screen for the simulated plant is located at /opt/rtcds/caltech/c1/medm/c1spx/master/C1SUP_ETMX.adl.  These are currently a work in progress.

  4379   Fri Mar 4 18:06:34 2011 kiwamuUpdateGreen Lockingnoise budget : differential noise

Here I explain how I estimate the contribution from the differential noise shown in the plot on my last entry (#4376) .

 

(background)

 According to the measurement done about a week ago, there is a broadband noise in the green beatnote path when both Green and IR are locked to the X arm.

The noise can be found on the first plot on this entry (#4352) drawn in purple. We call it differential noise.

However, remember, the thing we care is the noise appearing in the IR PDH port when the ALS standard configuration is applied (i.e. taking the beatnote and feeding it back to ETMX).

So we have to somehow convert the noise to that in terms of the ALS configuration.

In the ALS configuration, since the loop topology is slightly different from that when the differential noise was measured, we have to apply a transfer function to properly estimate the contribution.

 

(How to estimate)

 It's not so difficult to calculate the contribution from the differential noise under some reasonable assumptions.

Let us assume that the MC servo and the end PDH servo have a higher UGF than the ALS, and assume their gains are sufficiently big.

Then those assumptions allow us to simplify the control loop to like the diagram below:

servos.png

 Since we saw the differential noise from the beatnote path, I inject the noise after the frequency comparison in this model.

Eventually the noise is going to propagate to the f_IR_PDH port by multiplying by G/(1+G), where G is the open loop transfer function of the ALS.

The plot below shows the open loop transfer function which I used and the resultant G/(1+G).
 

open_loop_TF.png

In the curve of G/(1+G), you can see there is a broad bump with the gain of more than 1,  approximately from 20 Hz to 60 Hz.

Because of this bump, the resultant contribution from the differential noise at this region is now prominent as shown in the plot on the last entry (#4376).

Quote: #4376

I made a noise budget for the ALS noise measurement that I did a week ago (see #4352).

I am going to post some details about this plot later

 

  4378   Fri Mar 4 13:25:04 2011 ZachUpdateelogrestarted

with script

  4377   Fri Mar 4 09:47:46 2011 SureshUpdateElectronicsHarmonic distortion calculations for RF distribution box

 
Fast work indeed! It would be nice if we could have the following details filled in as well
a) A short title and caption for the table, saying what we are measuring
b) the units in which this physical quantity is being measured.

It is good to keep in mind that people from other parts of the group, who are not directly involved in this work, may also read this elog.

  4376   Fri Mar 4 03:31:35 2011 kiwamuUpdateGreen LockingA first noise budget

I made a noise budget for the ALS noise measurement that I did a week ago (see #4352).

I am going to post some details about this plot later because I am now too sleepy.

noise_budget.png

  4375   Thu Mar 3 20:30:03 2011 ranaSummaryPSLPMC Sweeps @ different input power levels to measure the Finesse

Its been well noted in the past that sweeping the PMC at high power leads to a distortion of the transmitted power curve. The explanation for this was coating absorption and thermo-elastic deformation of the front face of the mirrors.

Today, I did several sweeps of the PMC. I turned off its servo and tuned its PZT so that it was nearly resonating. Then I drove the NPRO via the HV driver (gain=15) with 0-150 V (its 1.1 MHz/V) to measure the PMC transmitted light. I adjusted the NPRO pump diode current from 2A on down to see if the curves have a power dependent width.

In the picasa web slideshow:

There are 3 significant differences between this measurement and the one by John linked above: its a new PMC (Rick says its the cleanest one around), the sweep is faster - since I'm using a scope instead of the ADC I feel free to drive the thing by ~70 MHz in one cycle. In principle, we could go faster, but I don't want to get into the region where we excite the PZT resonance. Doing ~100 MHz in ~30 ms should be OK. I think it may be that going this fast avoids some of the thermal distortion problems that John and others have seen in the past. On the next iteration, we should increase the modulation index for the 35.5 MHz sidebands so as to get a higher precision calibration of the sweep's range.

By eye I find that the FWHM from image #4 is 11 ms long. That corresponds to 300 mV on the input to the HV box and 15 V on the PZT and ~16.5 MHz of frequency shift. I think we expect a number more like 4-5 MHz; measurement suspicious.

  4374   Thu Mar 3 18:32:33 2011 Larisa ThorneUpdateElectronicsHarmonic distortion calculations for RF distribution box

Finished calculations for harmonic distortion at each of the 10 outputs of the RF distribution box. The diagram can be found on Suresh's post  http://nodus.ligo.caltech.edu:8080/40m/4342

 

THD calculation consisted of gather data on the dBm at harmonics of the fundamental frequency. These dBm values were converted into units of power and plugged into the appropriate THD equation pulled from Wikipedia: 
\mbox{THD} = \frac{P_2 + P_3 + P_4 + \cdots + P_\infty}{P_1} = \frac{\displaystyle\sum_{n=2}^\infty P_n}{P_1}

On the table, the number 1-6 correspond to the harmonic number of the input frequency used. For example, the first five PD's listed used an 11MHz source, while the second set of five PD's listed used a 55MHz source. Values listed under certain harmonics are dBm measurements at the corresponding frequency. The P-subscript values are essentially the dBm measurements converted to units of power (Watts) for ease of calculation in the equation above. THD is then calculated using these power units; I have converted the ratios to percentages.

 

It should be noted that as with all THD calculations, the more data points collected, the more precise the THD % will be.

By the way, the outputs on the physical RF distribution box for REFL165 and AS165 are actually labeled as REFL166 and AS166. 

Attachment 1: RFdistribboxcalcs.pdf
RFdistribboxcalcs.pdf
  4373   Thu Mar 3 07:25:24 2011 kiwamuUpdateGreen Lockingscrewed up the end PDH box

 I somehow screwed up the PDH box at the X end station. 

Right now it's not working, so I am going to check and fix it today.

 

 In the last evening I found that one of the gain stages on the PDH box wasn't fully functional.

So I started investigating it and I though it was going to finish soon, but actually it wasn't so easy.

 

  The PDH box has several gain stages. So an input signal goes through a buffer, a filter, a boost and an output buffer stages sequentially.

The boost stage is supposed to have gain of 10, but I found it didn't have such gain.

In fact the gain was something like -30dB which is pretty small. Plus this boost stage was imposing an wired bump on the transfer function around 50 kHz.

I checked the voltages on some components around the boost stage and confirmed there were no strange voltage.

Then I suspected that the op-amp : LF356 had been broken for some reason. So I replaced it by LT1792 to see if it fixes the issue.

Indeed it did make it functional. However after few minutes of the replacement, it went back to the same bad condition.

I have no idea about what was going on at that time. Anyway it needs more careful investigations.

 

  I temporarily put a jumper cable on the board to skip this stage, but now the PDH lock is not healthy at all.

  4372   Thu Mar 3 00:12:52 2011 kiwamuUpdateGreen Lockingplan
Tomorrow's tasks
  - Auto noise budget (Jamie)
  - Demodulation phase adjustment (Kate)
  - Auto alignment for green (Joe/Kiwamu)
  - ADC connection for the X end green REFL_DC ( )
  - remote local boost for the X end green ( )
  - TDS stuff (Joe)
  - check harmonic distortions on the RF distribution box (Larisa/Koji)
  - connect the X end mechanical shutter to c1auxex (Steve)
  4371   Wed Mar 2 22:57:57 2011 sureshSummaryGeneralStuff from LLO

Here is a partial list of stuff which is being packed at LLO to be shipped to CIT.  The electronics ckt boards are yet to be added to this list.  Will do that tomorrow.

 

 

Attachment 1: eLIGO_items_from_LLO_for_Caltech.xls
  4370   Wed Mar 2 22:04:22 2011 KevinUpdateElectronicsPOY Shot Noise Measurement

The previous measurement for the shot noise of POY had the dark noise at ~100 nV/rtHz. I redid the measurement and got 26 nV/rtHz for the dark noise. I think that when I made the previous measurement, the spectrum analyzer had automatically added some attenuation to the input that I failed to remove. This added attenuation raised the noise floor of the measurement making the dark noise of POY appear larger than it is.

The updated measurement can be found on the wiki at http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/POY.

  4369   Wed Mar 2 18:08:43 2011 AidanUpdateGreen LockingGhost beams on green

Kiwamu and I noticed that there is a ghost beam on the green beam going into the ETM. What we see is some interference fringes on the edge of the transmission of the green beam through the dichroic beam splitter (DCBS). If we look at the reflection from the dichroic beam splitter these are much more pronounced.

The spacing of the fringes (about 2 per 10mm) indicates an angle between the fields of around 0.1 mrad.

We were able to cause significant motion of the fringes by pushing on the knobs of the steering mirrors that steer the beam into the DCBS. A rough calculation of the derivative of optical path difference between the ghost and the primary beam as a function of input angle gives about 15 microns per mrad. What filtering the effect the arm cavity will have on the ghost beam is not immediately clear, but the numbers shouldn't be too difficult to determine.

 

Attachment 1: Ghost_Beam_at_ETM_DCBS.pdf
Ghost_Beam_at_ETM_DCBS.pdf
  4368   Wed Mar 2 17:19:58 2011 AidanConfigurationGreen LockingMoved PDH PD on end table

As previously noted, there are multiple beams coming back from the ETM surface onto the PDH PD on the end table. They are spread out in a vertical pattern. All the spots swing together (as the ETM moves?).

I moved the PDH Green PD on the end table so that it was further away from the Faraday and I added an iris in between the Faraday and the PD. Now only the principle reflection from the ETM is incident on the PD. See attached photos. In order to sneak past the neighbouring optics, I had to steer the beam down a bit, so the PD is now lower than it previously was.

Just FYI: the angle between the returning beams is about 5 or 6 mrad at the PD. Before the beams get to the PD they go through a telescope that de-magnifies the beam by about 5 or 6 times. This implies that the angle between adjacent returning beams from the ETM is around 1 mrad at the ETM.

This does make the position of the spot on the PD more susceptible to the alignment of the ETM - we should use a short focal length lens and image the ETM plane onto the PD.

 

First image - overview of table

Second image - the three returning beams immediately before the IRIS

Third image - a close up of the IRIS and PDH PD. 

 

 

 

Attachment 1: P1000223-a.jpg
P1000223-a.jpg
Attachment 2: P1000218.jpg
P1000218.jpg
Attachment 3: P1000224.jpg
P1000224.jpg
  4367   Wed Mar 2 16:51:53 2011 steveConfigurationGreen Lockingmech shutter in place at the south end

I moved old POX shutter from ITMY optical table to the south end. MEDM POX mechanical shutter screen is now closing the green beam  injection into the Y arm.

I kluged in a 40m long bnc cable that Alberto left on the floor for control. It is labelled POX-sht  This is a temporary set up.

  4366   Wed Mar 2 04:01:51 2011 KojiUpdateIOOMC servo improvement

[Koji / Rana]

- Since the MC servo had UGF up to ~20kHz and huge servo bump at 50kHz, we needed more phase between 20kHz to 100kHz.

- Today a phase compensation filter in a Pomona box has been inserted between the MC servo box and the FSS box.
  This is a passive filter with zero@14kHz and pole@140kHz. We obtain ~60deg at around 50kHz.

- After the insertion, the lock of the MC was achieved immediately. The overall gain as well as the PZT fast gain was tweaked
  such that the PC feedback is reduced down to 1~2.

- The OLTF has been measured.
  The insertion of the filter change increased the UGF to 130kHz even with "40:4kHz" and double super boost turned on.

  The phase margin is 54deg. Quite healthy.

- Rana modified the existed Auto Locker script.
  It is now continuously running on op340m!
  We made a couple of testsif it correctly relock the MC and it did. VERY COOL.

-----------------

NEXT STEPS
- Measure the PMC cavity pole
- Measure the circuit TF and try to shave off the phase lag.
- Measure the PZT resonance of the NPRO and put notch in the PZT path
- Increase the UGF / measure the openloop TF

Attachment 1: IMG_3904.jpg
IMG_3904.jpg
Attachment 2: MC_OLTF.pdf
MC_OLTF.pdf
  4365   Tue Mar 1 08:42:18 2011 AidanUpdateelogRestarted the elog this morning

 The elog was dead this morning. I reanimated it. It is now undead.

Attachment 1: Zombie.gif
Zombie.gif
  4364   Mon Feb 28 11:22:40 2011 josephbSummaryGeneralto do list
Quote:

- Where is the CDS TO DO ==> Joe

CDS To Do:

1) Get ETMY working - figure out why signals are not getting past the AI board (D000186) to the coils.

2) Get TDS and command line AWG stuff working

3) Get c1ass and new c1lsc (with Koji) fully integrated with the rest of the system.

4) Get CDS software instructions up to date and well organized.

5) Redo cabling  and generally make it a permanent installation instead of hack job:

   a) Measure cable lengths, check connectors, wire with good routes and ensure strain relief.  Ensure proper labeling

   b) Get correct length fiber for c1sus RFM and timing.

   c) Fix up final BO adapter box and DAC boxes.

   d) Make boxes for the AA filter adapters which are currently just hanging.

   e) Get two "faceplates" for the cards in the back of the ETMY IO chassis so they can screwed down properly.

   f) Remove and properly store old, unused cables, boards, and anything else.

6) Create new documentation detailing the current 40m setup, both DCC documents and interactive wiki.

7)  Setup an Ubuntu work station using Keith's wiki instructions

 

Simulated Plant To Do:

1)Create simulated plant to interface with current end mass controls (say scx).

2) Create proper filters for pendulum and noise generation, test suspension.

3) Propagate to all other suspensions.

4) Working on simulated IFO plant to connect to LSC.  Create filters for near locked (assume initial green control perhaps) state.

5) Test LSC controls on simulated IFO.

6) Fix c code so there's seamless switching between simulated and real controls.

CDS Status:

MC damp dataviewer diaggui AWG c1lsc c1ioo c1sus c1iscex c1iscey RFM The Dolphins Sim.Plant Frame builder TDS Cabling
                             
  4363   Sun Feb 27 13:09:56 2011 ranaUpdateGreen Lockingsidebands on beatnote

When Koji and I were massaging the MC, we noticed that the oscillations were at 48.5 kHz. They were pretty huge and are probably what you're seeing on the beat. My guess is that they are the PZT resonances of the PSL 2W NPRO; we need to put a notch in the FSS box - it still has the notch from the old NPRO.

  4362   Sun Feb 27 09:43:59 2011 AidanUpdateGreen Lockingsidebands on beatnote

Can we set up a fiber-PD on the end table to look at the beat between the "end laser IR beam" and the "PSL IR beam fiber-transmitted end beam"? 

We should see the same thing on that PD that we see on the green PD (plus any fiber noise and I'm not really sure how much that'll be off the top of my head). If we unlock the lasers from the arm cavity then the free-running noise of the lasers wrt to each other will probably swamp the 50kHz and 150kHz signals. Maybe we could lock the end laser to the free-running PSL by demodulating the beat note signal from the fiber-PD and then we could look for the extra sidebands in the IN-LOOP signal. Then we could progressively lock the PSL to the MC and arm cavity and see if the sidebands appear on the fiber-PD at some point in that process. 

It's possible that the 216kHz drive of the PZT on the Innolight is somehow driving up some sub-harmonics in the crystal. I think this is unlikely though: if you look at Mott's measurements of the Innolight PZT response, there are no significant PM resonances at 50 or 150kHz.

 

Quote:

Other than the +/-216 kHz sidebands, we can see some funny peaks at +/- 50 kHz and +/-150 kHz.

 

Quote: #4351 by Aidan

So, on the vertex PD, the power of the 80MHz +/-200kHz sidebands should be around sqrt(0.15)*0.05 = 0.02 = 2% of the 80MHz beatnote.

Once we get the green and IR locked to the arm again, we're going to look for the sidebands around the beatnote.

 

  4361   Sat Feb 26 02:33:16 2011 kiwamuUpdateGreen Lockingsidebands on beatnote

The power ratio of the beatnote signal vs. the 216kHz sideband has been measured.

The measured ratio was -55 dB, which is smaller by about 20 dB than Aidan's estimation.

To confirm this fact we should check the modulation depth of the end PDH somehow.

 

The below is a picture showing the sidebands around the beatnote locked at 66.45 MHz.

Other than the +/-216 kHz sidebands, we can see some funny peaks at +/- 50 kHz and +/-150 kHz

I wonder if they come from the servo oscillation of the MC servo whose UGF is at 24 kHz.  We can check it by unlocking the MC.

beat_note.png

Quote: #4351 by Aidan

So, on the vertex PD, the power of the 80MHz +/-200kHz sidebands should be around sqrt(0.15)*0.05 = 0.02 = 2% of the 80MHz beatnote.

Once we get the green and IR locked to the arm again, we're going to look for the sidebands around the beatnote.

  4360   Sat Feb 26 00:25:38 2011 KojiUpdateIOOMC servo improvement

[Rana / Koji]

The MC servo loop has been investigated as the MC servo was not an ideal state.

With the improved situation by us, the attached setting is used for the MC and the FSS.
The current UGF is 24kHz with phase margin is ~15deg, which is unbearably small.
We need to change the phase compensation in the FSS box some time in the next week.


- We found the PD has plenty of 29.5MHz signal in in-lock state. This was fixed by reducing the LO power and the modulation depth.

- The LO power for the MC demodulator was ~6dBm. As this was too high for the demodulator, we have reduced it down to 2dBm
by changing attenuator to 12dB (at 6 oclock of the dial) on the AM stabilization box.

- The RF power on the MC PD was still too high. The PD mush have been saturated. So the modulation slider for 29.5MHz was moved
from 0.0 to 5.0. This reduced the 29.5MHz component. (But eventually Koji restored the modulation depth after the servo shape has been modified.)

- The openloop gain of the loop has been measured using EXC A/TEST1/TEST2. The UGF was ~5kHz with the phase mergin of ~10deg. 

- This quite low phase margin is caused by the fact that the loop has f^-2 shape at around 4k-100kHz. The reference cavity has
the cavity pole of 40kHz or so while the IMC has the pole of ~4kHz. Basically we need phase lead at  around 10-100kHz.

- We decided to turn off (disable) 40:4000 boost of the MC servo to earn some phase. Then MC did not lock. This is because the LF gain was not enough.
So put Kevin's pomona box in the FAST PZT path (1.6:40). By this operation we obtain ~75deg (max) at 560Hz, ~35deg at 5kHz, ~20deg at 10kHz.

- In this setup the UGF is 24kHz. Still the phase margin is ~15kHz. This phase lag might be cause by 1)  the MC servo circut 2) PMC cavity pole

NEXT STEP

- Put/modify phase lead in the FSS box.
- Measure the PMC cavity pole
- Measure and put notch in the PZT path
- Increase the UGF / measure the openloop TF

Attachment 1: fss_servo.png
fss_servo.png
Attachment 2: mc_servo.png
mc_servo.png
  4359   Fri Feb 25 14:50:16 2011 KojiSummaryGeneralto do list

- Put priority on the list

- Put names on the items

- Where is the CDS TO DO ==> Joe

-

- Remote disconnection of the greeen PDH 

- What is the situation of the PD DC for the LSC PDs?

- SUS Satelite box Resister replacement ==> Jamie

- IMC mode matching ==> Jamie/Larisa 

- Mechanical shutters everywhere

- SRM OPLEV Connection

- MC OAF

- Better LSC whitening boards

- DAFI 

Quote:
 Anything else ?

  4358   Fri Feb 25 14:35:06 2011 Larisa ThorneUpdateElectronicsTotal harmonic distortion results for +7dBm mixer

 This experiment deals with measuring the total harmonic distortion (THD) contribution of mixers in a circuit.

(a circuit diagram is attached) where:

Mixer: ZFM-3-S+ at +7dBm

Attenuator: VAT-7+, at +7dB

Low-pass filter: SLP-1.9+, which is set to DC-1.9MHz

 

The total harmonic distortion can be calculated by the equation: 


\mbox{THD} =  \frac{V_2^2 + V_3^2 + V_4^2 + \cdots + V_\infty^2}{V_1^2}

where Vn represents the voltage of the signal at a certain harmonic n.

 

In this experiment, only the voltages of the first three harmonics were measured, with the first harmonic at 400Hz, the second at 800Hz, and the third at 1.2kHz. The corresponding voltages were read off the spectrum analyzer after it had time averaged 16 measurements. (picture of the general shape of the spectrum analyzer output is attached)


(results for this mixer's particular configuration are on the pdf attached)

 

There really isn't that much correlation between the modulations and the resulting THD.

We won't know how good these numbers are until more experiments on other mixers are done, so they can be compared. Since the rest of the mixers are relatively high levels (+17dBm, +23dBm in comparison to this experiment's +7dBm), an RF amplifier will need to be hooked up first to do any further measurements.

 

 

 

 

 

Attachment 1: THDcircuit.jpg
THDcircuit.jpg
Attachment 2: Photo_on_2011-01-17_at_12.25.png
Photo_on_2011-01-17_at_12.25.png
Attachment 3: THDwithoutamp.pdf
THDwithoutamp.pdf
  4357   Fri Feb 25 13:28:14 2011 kiwamuSummaryGeneralto do list
 Because it maybe useful to share this kind of information among us, I just put a to-do-list which has been continuously noted in my laptop.
 I guess those items can be applied for the coming two months (i.e. March and April). Anything else ?

-- Alignment

     - C1ASS

           * finalize the model and make user friendly scripts and medm screens
           * adjustment of phase rotations
           * cavity auto alignment for green

     - MC WFS 

          * put sensing matrices
          * whitening ?

    - OPLEVs

        * binary output for OPLEV whitening
        * calibration of OPLEVs  [rad/count]
        * characterization of OPLEV servos
        * bigger trans impedance gains to have reasonable readouts


-- Optimization of Suspended Optics

  - diagonalizations

         * input matrices
         * f2p with higher precision

  - damping control

        * fix ETMY damping
         * Bounce Roll notches
         * Q adjustment for damping

  - upgrading of electronics

        * bigger trans impedance gain for OSEMs
        * widen the voltage range of AA filter boards
   

   - weekly check

       * a routine script for measuring free swinging spectra


-- Input Optics

   - PMC

         * mode matching
         * epics LO HI values

   - FSS and ISS

        * recover FSS
        * make ISS working

   - EOM

       * AM minimization
       * triple resonant box

   - doubling and RFPD for green

        * mode matching to doubling crystal
        * connect RFPD_DC_MON to ADC
        * string +/-150 V and +/-15power cables from 1X1 rack to RFPDs
        * visibility check and loss investigations for the beat RFPD
        * rearrange RF amplifiers (ZLN series) for the RFPD
        * realgin Jenne's DCPD

 

-- Length Sensing and Control

   - digital system and electronics

       * characterization of RFPDs ==> SUresh/UG
       * installation of RF generation box and distribution box ==> Suresh/U
       * new LSC model and start making useful scripts (csh ? perl ? python ?)
       * binary outputs for PD whitening
       * make item lists for ordering (?)
       * draw cool diagrams for RF cable distribution and map of LSC rack.


-- Green Locking

    - X end station

           * eliminate undesired multiple spots on RFPD
           *  connect REFLPD_DCMON to ADC
           *  remote local boost
           * demodulation phase adjustment
           *  look for a high voltage amp. (bipoler)
           * installation of a mechanical shutter
           *  ETMX_TRANS CCD camera
           * analog low pass filter for temperature control

    - Y arm green locking ==> Suresh/Bryan

        * item lists
        * preparation of base mounts
        * mode matching estimation
        * mode measurement of input beam from Lightwave
        * temeperature scanning for beat location finding with IR beam
        * Installation
       * modification of PDH box

    - digital control systems

           * user friendly medm screens
           * apply proper filters for AC-DC whitened signal blending
           * add MC2 feedback path

      - noise budget

          * in-loop and out-of-loop evaluation
          * shot noise  
          * RFPD noise modeling
          * how intensity noise couples through MFD
          * electrical noise
          * frequency noise contribution from end laser and PSL
          * calibration of arm PDH signal

   - cavity scan and handing off

         * optimization of open loop transfer function for ALS
         * auto scripts   

-- misc.

   - CCD camera
        * color filters to separate IR and green (?)
   - lab laptops
        * a laptop for each end station (?)
    - dichroic TT (?)
        * large wegded and AR coated for 532nm (?)
    - epics for RS232C      
        * RS232C for doubling oven temperature

 

  4356   Fri Feb 25 10:11:56 2011 steveUpdateSAFETYhow not to

The beam of IR for doubling  is clipping on bnc cable to green beam transmitted pd.

Attachment 1: Presentation1.jpg
Presentation1.jpg
  4355   Fri Feb 25 01:48:54 2011 valeraUpdateASCmc auto alignment status

 I made several scripts to handle the mcass configuration and sensing measurements:

- The scripts and data are in the scripts/ASS directory

- The mcassUp script restores the settings for the digital lockins: oscillator gains, phases, and filters. The MC mirrors are modulated in pitch at 10, 11, 12 Hz and in yaw at 10.5, 11.5, and 12.5 Hz. The attached plot shows the comb of modulation frequencies in the MCL spectrum.

- The mcassOn and mcassOff scripts turn on and off the dither lines by ramping up and down the SUS-MC1_ASCPIT etc gains

- The senseMCdecenter script measures the response of the MCL demodulated signals to the decentering of the beam on the optics by imbalancing the coil gains by 10% which corresponds to the shift of the optic rotation point relative to the beam by 2.65 mm (75mm diameter optic) and allows calibration of the demodulated signals in mm of decentering. The order of the steps was MC1,2,3 pitch and MC1,2,3 yaw. The output of the script can be redirected to the file and analyzed in matlab. The attached plot shows the results. The plot was made using the sensemcass.m script in the same directory.

- The senseMCmirror script measures the response of the MCL demodulated signals to the mirror offsets (SUS-MC1_ASCPIT etc filter banks). The result is shown below (the sensemcass.m script makes this plot as well). There is some coupling between pitch and yaw drives so the MC coils can use some balancing - currently all gains are unity.

- The senseMCdofs scripts measures the response to the DOF excitation but I have not got to it yet.

- The next step is to invert the sensing matrix and try to center the beams on the mirrors by feeding back to optics. Note that the MC1/MC3 pitch differential and yaw common dofs are expected to have much smaller response than the other two dofs due to geometry of this tree mirror cavity. We should try to build this into the inversion.

Attachment 1: mcditherlines.pdf
mcditherlines.pdf
Attachment 2: mcdecenter.pdf
mcdecenter.pdf
Attachment 3: mcmirror.pdf
mcmirror.pdf
  4354   Thu Feb 24 21:46:30 2011 kiwamuUpdateGreen Lockinginstalled a summing box

In this past weekend I replaced a summing amplifier for the end green PDH locking by a home-made summing circuit box in order to increase the control range.

It's been working well so far.

However due to this circuit box, the demodulation phase of the PDH locking is now somewhat different from the past, so we have to readjust it at some point.

 

(background)

    At the X end station, the voltage going to the NPRO PZT had been limited up +/- 4 V because of the summing amplifier : SR560.

Therefore the laser was following the cavity motion only up to ~ +/- 4 MHz, which is not wide enough. (it's okay for night time)

So we decided to put a passive circuit instead of SR560 to have a wider range.

 

(summing box)

   We made a passive summing circuit and put it into a Pomona box.

The circuit diagram is shown below. Note that we assume the capacitance of the 1W Innolight has the same capacitance as that of the PSL Innolight (see #3640).

summing_box.png

The feedback signal from a PDH box goes into the feedback input of the circuit.

Then the signal will be low passed with the corner frequency of 200 kHz because of the combination of RC (where R is 681 Ohm and C is capacitance of the PZT).

Because of this low pass filter, we don't drive the PZT unnecessarily at high frequency.

On the other hand the modulation signal from a function generator goes into the other input and will be high passed by 50 pF mica capacitor with the corner frequency of 200 kHz.

This high pass filter will cut off noise coming from the function generator at low frequency.

In addition to it, the 50 pF capacitor gives a sufficient amount of attenuation for the modulation because we don't want have too big modulation depth.

 

Here is a plot for the expected transfer functions.

You can see that the modulation transfer function (blue curve) has non-zero phase at 216 kHz, which is our modulation frequency.

transfer_func.png
 

 

  4353   Thu Feb 24 19:59:25 2011 kiwamuUpdateGreen Lockingwhitening filter for ALS

I forgot to mention about the whitening filter for the ALS digital control system.

As usual I used a whitening filter to have a good SNR against ADC noise, but this time our whitening scheme is little bit different from the usual our systems.

I used two ADC channels for one signal and put a digital summing point  and digital filters to keep good SNR over the frequency range of interest.

It's been working fine but it's still primitive, so I will study more about how to optimize this scheme.


ACDC.png

     The diagram above shows our scheme for the signal whitening.

Basically the SNR at DC is bad when we use only a whitening filter as shown on the bottom part of the diagram because the signal is quite tiny at DC.

On the other hand if we take raw signal into ADC as 'DC path'  shown above, the SNR is better at DC but not good at intermediate frequencies (30 mHz - 1kHz).

So the idea to keep the good SNR over the frequency range of interest is to combine these 'DC path' and 'AC path' in a clever way.

     In our case, the 'DC path' signal is not as good as the 'AC path' signal above 30 mHz, so we cut off those high frequency signals by using a digital low pass filter.

In addition to it, I put a gain of 1000 in order to match the relative gain difference between 'DC path' and 'AC path'.

Then the resultant signal after the summing point keeps the good SNR with a flat transfer function up to 1 kHz. 

Quote:

Two different measurement have been performed for a test of the green locking last night.

Everything is getting better. yes. yes.

ELOG V3.1.3-